gamerboy:

Mya, Jean, and Cody shared a pie. May ate 5/12 of the pie, Jean ate 1/12 of the pie, and Cody ate 1/2 of the pie. • Part A: What fraction represents the total amount of the pie that Mya and Cody ate? Show your work and explain how you determined your answer. • Part B: Mya said she ate a larger fraction of the pie than Cody did. Explain whether Mya is correct.

gamerboy:

@Mercury

kittybasil:

Okay, we have two people who ate "twelfths" (based on the fractions) of the pie. That's Mya (or is it May?) and Jean. Cody ate half of the pie. This isn't a twelfth, but we can convert the fraction to that. What's half of 12? Six -- therefore, Cody has also eaten 6/12 of the pie.

kittybasil:

So now we have:\[\text{May/Mya ate }\frac{5}{12}\]\[\text{Jean ate }\frac{1}{12}\]\[\text{Cody ate }\frac{6}{12}\]Now these are all the same kind of fraction, so you can add them.

kittybasil:

Are you with me so far? @gamerboy Let me know if you are stuck or confused anywhere.

gamerboy:

yes im with you. i have to add the fractions together?

kittybasil:

Yes :-) For part A, you will add the fractions for Mya (?) and Cody together. This means you add:\[\frac{5}{12}+\frac{6}{12}=?\]Don't forget to show the part where you converted Cody's fraction of the pie to this fraction format. For part B, you have to prove whether Mya ate more than Cody. Which means:\[\frac{5}{12}\stackrel{?}{\gt}\frac{6}{12}\]Is the statement above true?

gamerboy:

5/12 + 6/12 = 11/12

gamerboy:

the above statement is false?

kittybasil:

For part B, Mya's claim is false. Because\[\frac{5}{12}\cancel{\gt}\frac{6}{12}\]

kittybasil:

Sorry -- what I meant to say is you're correct, just clarifying the details.

gamerboy:

noo ur fine thank you so much for helping!

kittybasil:

No problem, good luck on your studies!\[\Huge{\ddot{\smile}}\]

gamerboy:

oh btw im sorry for not telling you that the name is actually Mya. that was just a little mistake i made there